OpenStudy (anonymous):

5+4z=17. i completely forget how to do these

OpenStudy (amistre64):

you keep everything balanced by performing the same operations to both sides

OpenStudy (amistre64):

in this case; -5 and divide off the 4

OpenStudy (anonymous):

17-5=4z 4z=12 z=12/4 z=3